OpenStudy (anonymous):

find the first derivative of y = sin(2x) sin(3x) with solution

OpenStudy (john_es):

\[y'=(\sin(2x))'\sin(3x)+\sin(2x)(\sin(3x))'=2\cos(2x)\sin(3x)+3\sin(2x)\cos(3x)\]